A certain cable car is San Francisco can stop in 10 s when traveling at maximum speed. On one occasion, the driver sees a dogA certain cable car is San Francisco can stop in 10 s when traveling at maximum speed. On one occasion, the driver sees a dog a distance d m in front of the car and slams on the brakes instantly. The car reaches the dog 7.92 s later, and the dog jumps off the track just in time. If the car travels 3.83 m beyond the position of the dog before coming to a stop, how far was the car from the dog? [Hint: You will need three equations.] m
Answer #1

lets say the initial speed is n and the relardation of the car is (-a). Time to stop Vaatat (v=0 when stops) - O = u-at - a+7.928 S = ut t at² (+= 7.925 ut 17.7m/s Q- - 1.77m18?) + 1 x (-1.77)x(4.9% = 17.7x7.92 u to it.tms. Sr 84.67m Cans)